OpenStudy (anonymous):

(2.1X10^-3)(6X10^-2)/4.2X10-4 please help thank you

OpenStudy (saifoo.khan):

Convert things into same exponents of 10.

OpenStudy (saifoo.khan):

\[\Large \frac{(2.1 \times 10^{-3} )(60 \times 10^{-3})}{0.42 \times 10^{-3}}\]

OpenStudy (anonymous):

wait yy would i converet can u please explain it better and just write the answer

OpenStudy (saifoo.khan):

Well, we convert them so we bring them all on the same track and then solve them.

OpenStudy (anonymous):

can u solve it

OpenStudy (saifoo.khan):

Now take commons, \[\Large \frac{(2.1 * 60) \times 10^-3}{0.42 \times 10^{-3}}\]\[\Large \frac{2.1 \times 60}{0.42} \times \frac{10^{-3}}{10^{-3}}\]

OpenStudy (anonymous):

wow lol thank u very much

OpenStudy (saifoo.khan):

So wait i did a mistake.

OpenStudy (saifoo.khan):

Oh*

OpenStudy (saifoo.khan):

\[\Large \frac{(2.1 * 60) \times (10^{-3} * 10^{-3})}{0.42 \times 10^{-3}}\]It will be something like this.

OpenStudy (anonymous):

kkk thanks

OpenStudy (saifoo.khan):

So,